University of Vermont

UVM Blogs

fcs

  • 28 October 2014: Tuesday

    fcs on Frank's Activity Log
    LDAP: Cycle for new schema for that two year old project. Deal with issues in last night’s update: Blank line – Hello Account Services… Wrong surname with one datum match: Appears that OSP put the […]
    Posted yesterday.
  • 27 October 2014: Monday

    fcs on Frank's Activity Log
    LDAP: Doh!  No issues over the weekend.  Why can’t NetWorker be more like LDAP??? 1,430 BCA only affiliations removed after the change went into production last week.  YAY! List of removals sent to […]
    Posted 2 days ago.
  • 24 October 2014: Friday

    fcs on Frank's Activity Log
    LDAP: Issues: Two former students from Banner – did the dance. Added monitoring for the temporary insecure idiotic ldap server that Oracle will be allowed to use until they can actually do something […]
    Posted 5 days ago.
  • 23 October 2014: Thursday

    fcs on Frank's Activity Log
    LDAP: One issue last night – a “modifying the WRONG entry” error: Employee record gained an SSN and now it matches a former student LDAP entry.  Merged the Banner info into the new HR entry (because the HR […]
    Posted 6 days ago.
  • 22 October 2014: Wednesday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s update. POODLE recycles. POODLE (CVE-2014-3566) recycles – take two. Network Services feed – updated. Backups: CoM/IS clones have about 8.5TB working and […]
    Posted 1 weeks ago.
  • 21 October 2014: Tuesday

    fcs on Frank's Activity Log
    LDAP: No Issues last night. I think I have found the solution… TLSProtocolMin 3.1 Backups: Tape clones finished, the rest of them run on. MED23 B continues to perform the full from the […]
    Posted 2 weeks ago.
  • 20 October 2014: Monday

    fcs on Frank's Activity Log
    LDAP: One issue over the weekend. A Visiting Graduate Student who didn’t match up.  Found the match and merged. Backups: Opened SR 66570742 to deal with the fact that NetWorker is not properly […]
    Posted 2 weeks ago.
  • 17 October 2014: Friday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s update. Backups: Disk clones have 13T left (at 7:41AM) NetWorker frozen in place.  Apparently, the “only use X% of a disk” option is too stupid to be used.  NetWorker […]
    Posted 2 weeks ago.
  • 16 October 2014: Thursday

    fcs on Frank's Activity Log
    LDAP: Two one-item match, wrong surname issues in last night’s update: Both PeopleSoft: One is a former student – asking for confirmation of details.  HR confirms, merged and account created. One is […]
    Posted 2 weeks ago.
  • 15 October 2014: Wednesday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s update. Today is Automated Account Maintenance day for October.  It ran with four reported errors (trying to add email bounce controls that already existed). Cleaned up an […]
    Posted 2 weeks ago.
  • 14 October 2014: Tuesday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s update. Backups: Clones:  CoM/IS is almost done, Disk has about 36TB left to go. DataDomain: Is working away on moving the data to tape.  I hope it frees up enough […]
    Posted 3 weeks ago.
  • 13 October 2014: Monday

    fcs on Frank's Activity Log
    LDAP: Two issues over the weekend: student employees that have not ssn’s yet.  Found one and merged it, asking about the other.  The second one is not a student, but a “visiting scholar” who was entered […]
    Posted 3 weeks ago.
  • 10 October 2014: Friday

    fcs on Frank's Activity Log
    LDAP: VPN cleanup per request. Creating a retiree. No issues in the last two nights updates. Backups: email backups are taking too long. CoM/IS clones finished. Disk clones aborted to free up […]
    Posted 3 weeks ago.
  • 8 October 2014: Wednesday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s update. Merged two entries that are determined to be the same person. Backups: EMC NW SR # 66189270: Oops, that code I wrote yesterday didn’t run (even though -cw said […]
    Posted 3 weeks ago.
  • 7 October 2014: Tuesday

    fcs on Frank's Activity Log
    LDAP: No issues in last night’s update. Backups: Continuing to relabel the tapes that came back from the vaults yesterday. EMC NW SR # 66189270 – writing code to stage one SSID at a time to provide […]
    Posted 4 weeks ago.
  • 6 October 2014: Monday

    fcs on Frank's Activity Log
    LDAP: Issues.  There are supposed to be issues.  Why are there not any issues? Backups: Tape Drive with a code 6 – cleaning it cleared the issue – was it just dirty? EMC DD SR # 66027568: Wants logs […]
    Posted 4 weeks ago.
  • 3 October 2014: Friday

    fcs on Frank's Activity Log
    Been out sick Tuesday and Wednesday this week, Thursday was working from home for just over half the day to try to get caught up on email and a few of the issues (Data Domain & NetWorker & SecurID). LDAP: No […]
    Posted 4 weeks ago.
  • 29 September 2014: Monday

    fcs on Frank's Activity Log
    LDAP: No Issues with the nightly updates over the weekend. Backups: Email backups continue to run to tape.  There was tape drive contention that delayed the start of the backups by almost 12 hours. I […]
    Posted 5 weeks ago.
  • 26 September 2014: Friday

    fcs on Frank's Activity Log
    LDAP: One Issue in nightly update: Former Student: did the dance. CoM/IS retrieval stuff. Backups: Data Domain at 0.0 free again. Compression has been reduced since we changed on 8/6 […]
    Posted last month.
  • 25 September 2014: Thursday

    fcs on Frank's Activity Log
    LDAP: No issues in the nightly update. Backups: The Data Domain ran out of space.  Backups are halted. SR 65904996 – EMC insisting that something unexpected has happened and that we really need to […]
    Posted last month.